I've looked over this word problem several times and I can't figure out what to do, here is the problem:
North Carolina's Native-American population is 22,000 greater than New York's. New York's Native'American population is 187,000 less than Oklahoma's. If the total population of all three is 437,000, find each state's Native- American population.
Please help me, I don't know what to do. Thank you very much!

Answer by queenofit(120)

let x equal oklahoma's population
let x - 187,000 equal new york population
let x- 187,000+ 22,000 equal North carolina
(x-187,000)+ (x-187,000 + 22,000)+ x = 437,000
3x - 352,000 = 437,000
3x=352,000+437,000
3x=789,000
x=263,000 ok population
plug x answer into other problems
263,000-187,00=76,000 New yorks population
263,000-187,000+22,000 = 98,000 North Carolina population
check answer
263,000 + 76,000 + 98,000 = 437,000